OpenStudy (christos):

For the functions f(x) = 9x - 2 and g(x) = 4x - 6 find FoG and state its domain

OpenStudy (anonymous):

\[f o g(x) = f(g(x))\] and its read "f of of x"

OpenStudy (anonymous):

f(x) = 9x - 2 g(x) = 4x - 6 f(g(x)) substitute g(x) into f(x) f(g(x)) = 9(4x - 6) - 2 f(g(x)) = 36x - 54 - 2 f(g(x)) = 36x - 56

OpenStudy (anonymous):